# A Survey on Error-bounded Lossy Compression for Scientific Datasets

### v2

@@ 6. Customized Compressors for Specific Applications or Use Cases
In this section we carefully survey the error-bounded lossy compressors that were tailored for specific applications/use cases.
@@ 6.1 Compression for Molecular Dynamics Simulations
Lossy compression has been widely considered a promising solution to reduce the data volume of MD simulations. For example, GROMACS (Hess et al., 2008), which is one of the leading MD simulation packages, has had its lossy format XTC built-in for decades. However, designing lossy compressors for MD simulations also presents unique challenges. First, the dominant type of data that needs to be stored in MD simulations is the particle trajectory, which is made of multiple frames (snapshots) of particle coordinates in a 3D space. Other data, including the particle velocities, forces, and system topology, either are not required in many cases or take much less storage than the trajectory. Compared with the structured mesh (regular multidimensional grid) that many other scientific applications use, the trajectory format is different: it stores discrete coordinates whereas the mesh stores continuous values, so that the existing error-bounded lossy compressors such as SZ and ZFP may not work effectively. Second, although the MD trajectory has a temporal dimension, it is not feasible to treat the data as time series to compress because of the requirement of random access (access data from randomly selected frames) for postanalysis, such that the compression needs to be done in batches, each containing only a limited number of frames.
Given these challenges, there is ongoing research into lossy compression methods tailored for MD simulations. The HRTC method (Huwald et al., 2016) employs a strategy that represents trajectories as piecewise linear segments, coupled with quantization that is controlled for errors and a representation using variable-length integers. The PMC approach (Dvořák et al., 2020) leverages the information about atomic bonds within molecules to forecast the positions of atoms in each frame; however, this technique does not apply to simulations involving nonbonded interactions. Omeltchenko et al. (Omeltchenko et al., 2000) proposed a spatial compressor for MD datasets that includes three steps: (1) converting all floating-point values (both position and velocity) to integer numbers, (2) building a uniform oct-tree index according to the space-filling curve of the position fields, and (3) sorting the particles based on R-indices using a radix-similar sorting method in each block and encoding the difference in adjacent indices by variable-length encoding. Tao et al. (Tao et al., 2017b) improved Omeltchenko et al.’s method by sorting the particles based on a partial-radix sorting algorithm while preserving the same compression ratios by using SZ to compress the reordered coordinates instead of directly using the R-index. Essential dynamics (Meyer et al., 2006) is a powerful analysis tool for identifying the nature and relative importance of the essential deformation modes of a macromolecule from MD samplings. ED offers lossy compression by adopting principal component analysis (PCA) over the full trajectories of all particles. By comparison, Kumar et al. (Kumar et al., 2013) combined PCA and discrete cosine transform (DCT) to compress the full trajectories of the total MD dataset. Such full-trajectory-based compression methods, however, may be impractical for many of today’s large-scale MD simulations. Zhao et al. (Zhao et al., 2022) proposed an SZ2-based compressor, MDZ, which is equipped with spatial-clustering-based prediction and two-level temporal prediction. MDZ achieves high compression ratios particularly on MD simulations focusing on crystalline materials or featuring a continuous temporal domain.
@@ 6.3 Compression for Quantum Circuit Simulations
Wu et al. (Wu and et al., 2019) designed a Schrödinger algorithm-based full state vector simulation pipeline that integrates compression. MPI is used to parallelize the matrix multiplication required to apply a gate to a state vector. Each rank stores a set of compressed blocks that together compose a component of the overall state vector. Blocks are decompressed two at a time to perform a partial state vector update, and the resulting state vector piece is compressed for later use. Wu et al. explored multiple compressors to compress the state vector data, including SZ2.1 (A), SZ2.1 with complex type support (B), XOR leading-zero reduction coupled with bit-plane truncation and zstd (C), and reshuffling of real and imaginary parts together before performing C (D). The results indicate that solutions C and D achieve the highest compression ratios (30–90 for error bounds in the range [1E-5,1E-1]) across the four configurations including FPZIP and ZFP. When running a 61-qubit Grover’s search algorithm with this compressor, the memory requirement drops from 32 exabytes to 768 terabytes using 4,096 nodes. In all, their method can raise the number of qubits for a simulation by 2 to 16.
@@ 6.6 Compression for Topological Data Analysis
Topological data analysis is essential in abstracting, summarizing, and understanding scientific data in various applications, ranging from cosmology and combustion to Earth simulations and AI. Topological feature descriptors, or simply topological descriptors, provide robust capabilities for capturing, summarizing, and comparing features in scientific data. Most lossy compressors cannot preserve topological features, thus not guaranteeing topology preservation in decompressed data. Inconsistency of topology in decompressed data could lead to misinterpretation and even wrong discoveries. Below is a review of the preservation of topological features in error-bounded lossy compression in two aspects: scalar field topology and vector field topology, both of which require customizations of existing lossy compressors.
@@ 6.7 Compression for Multi-resolution Data
Multi-resolution methods, such as the Adaptive Mesh Refinement (AMR) technique have been widely used in scientific simulations (Zhang et al., 2019a; Stone et al., 2020; Dubey and et al., 2022). AMR aims to reduce computational expenses while preserving the accuracy of simulation outcomes. Unlike traditional uniform mesh techniques that apply consistent resolution throughout the simulation space, AMR employs a dynamic approach. It selectively increases resolution in regions of interest, thereby optimizing computational resource usage and minimizing storage requirements. However, the space saved from using multi-resolution alone is often not enough. For instance, a multi-resolution dataset with $0.5\times 1024^{3}$ mesh points at the coarse level and $0.5\times 2048^{3}$ at the fine level could yield about 1 TB of data per snapshot. Consequently, conducting five simulations with 200 snapshots would require a total disk storage of 1 PB. Simulations used in Exascale scenarios can be even larger than that, using many thousands of points per axis (Fedeli and et al., 2022), making data size reduction a timely need.
While lossy compression has the potential to significantly reduce I/O and storage costs for multi-resolution data, a gap exists between compression and multi-resolution techniques especially because of the hierarchical nature of AMR data, where the entire dataset possesses varying resolutions/levels. Data at each level is sparse because it only covers a portion of the domain. Yet, current scientific compressors exclusively support the compression of non-sparse data with uniform resolution. Consequently, AMR data must be pre-processed before compression.
@@ 6.8 Compression for Seismic Imaging
Seismic imaging is a technique for determining the seismic properties of the Earth’s subsurface (Li and Qu, 2022). The technique is extensively utilized in earthquake imaging and resource exploration, including hydrocarbon and geothermal, by energy companies such as Saudi Aramco (Huang et al., 2023c). Among all the existing seismic imaging methods, reverse time migration (RTM) is a cutting-edge one since it can effectively analyze complex seismic structures (e.g., complex velocity focusing and steep (¿70) dips imaging), compared with traditional methods such as Kirchhoff and wave equation migration  (Farmer et al., 2009).
@@
Huang et al. (Huang et al., 2023c) proposed a hybrid lossy compression method called HyZ that combines blockwise regression (BR) and an ultra-fast prediction-based compressor SZx (Yu et al., 2022) to improve the performance of RTM overall execution. Evaluation on 3,600 snapshots of the Overthrust model shows HyZ achieves a compression ratio of 12.31x and compression/decompression speeds of 10.69 GB/s and 12.45 GB/s, respectively. Integration of HyZ into an industrial parallel RTM code improves overall performance by 6.29–6.60x over the execution without compression techniques, outperforming second-tier compressors like SZ and ZFP by up to 2.23x. HyZ also demonstrates higher fidelity than BR in preserving the visualization quality of single snapshots and the final stacking image.
@@ 6.9 Compression for X-ray Light Source Data
For serial crystallography, two major approaches can be combined: non-hit rejection and ROIBIN-SZ (Underwood et al., 2023b). Like ptychography, the data present as 2D float-encoded integer data over a time dimension. Unlike ptychography, however, the data are substantially noisier, and there are features called Bragg spots or peaks that are key to the analysis pipeline that need to be preserved more conservatively. Non-hit rejection is a technique that uses the number of peaks detected each frame to veto or reject capturing frames that contain few, if any, peaks. While non-hit rejection eliminates on average 50% (typically between 20% and 80%) of the data from an experiment, it alone is not enough to hit the compression ratio targets for these workflows. In order to achieve higher compression ratios, a method called ROBIN-SZ was developed. ROIBIN-SZ uses the peak information used to perform non-hit rejection to losslessly preserve rectangular regions around the Bragg peaks, while using aggressive $2\times 2$ binning followed by SZ3 compression on the background. Preserving the background is critical because the peak-finding process is not infallible. There could be false-negative peaks to preserve in the dataset in order for the analysis process to complete.
@@ 6.10 Compression for Data Transfer over WAN
Recently, error-bounded lossy compression techniques have also been used to improve the data transfer performance over the wide area network (WAN). Ocelet (Liu et al., 2023a)—a lossy-compression-based data transfer accelerator developed for the Globus platform—is a typical example. The Ocelet framework is composed of 8 components/modules: user interface, FuncX service (fun, [n. d.]), Globus service (glo, [n. d.]), parallel executor, MPI call module, error-bounded lossy compression module, data loader/writer, and lossy compression quality estimation module. The lossy compression quality estimation module is used to find a suitable error bound and compressor to conduct the compression. The data loader is used to load the data of multiple formats such as NetCDF (Rew and Davis, 1990), HDF5 (hdf, [n. d.]), and binary. The parallel executor is used to launch the compression/decompression work in a parallel job. Globus manages the data transfer. FuncX service deals with remote orchestration. The user interface offers a graphical interface that helps users submit the tasks easily.
@@ 6.11 Compression for Boosting Communication in HPC Clusters
Researchers have been actively investigating the application of lossy compression to boost the performance of communication in high-performance clusters, focusing on two primary categories: point-to-point and collective communication. Collective communication operations encompass a variety of types, which fall into two distinct subcategories, collective data movement and collective computation, based on their respective communication patterns.
@@ 6.12 Compression for Distributed Machine Learning & Federated Learning Systems
Recent years have witnessed the rapid evolution of deep learning models for getting high model accuracy, especially in the realm of large-scale models (Bommasani et al., 2022). Typical examples include large language foundation models (e.g., Palm and GPT-4 (Anil et al., 2023; OpenAI et al., 2023)), large-scale models in computer vision (e.g., VGGs, ResNets (Simonyan and Zisserman, 2015; He et al., 2016)), and life science (e.g., AlphaFolds (Jumper and et al., 2021)). Distributed systems, including the public/private cloud and HPC, provide strong support for training large-scale models based on large-volume training datasets to accelerate the training procedure (Li et al., [n. d.], 2013; Zhang et al., 2022; Zhang and Wang, 2021; sergeev2018horovods; Xu et al., 2024).
Communication between computing nodes during training has been recognized as the primary bottleneck in distributed training systems. During the training, gradients, model parameters, and activation data are transmitted across various nodes. As model sizes increase, the volume of this data escalates significantly, resulting in substantial communication overhead. To enhance training performance, it is essential to compress these data to reduce communication demands. The communication reduction approaches include: ❶ reducing communication rounds, ❷ communication and computing overlapping, and ❸ gradients and parameter compression. The gradient compression approaches mainly consist of gradient sparsification, gradient quantization, low rank, and error-bounded lossy compression.
@@
Error-Bounded Lossy Compressors An emerging area of research is using error-bounded lossy compressors such as SZ and ZFP to reduce the size of model updates for distributed systems. Early approaches such as DeepSZ (Jin et al., 2019) focused on providing a framework for compression and storage of general DNN model architectures. Their results showed that through compressing the weights of the model, they could achieve high compression ratios in those areas, leading to reported 50$\approx$ 50$\times$ 50 × compression on the weights. Other studies, such as the FedSZ framework (Wilkins et al., 2023), demonstrate how SZ-based compression can be integrated into federated learning (FL). By compressing local model updates before transmission, FedSZ effectively reduces the bandwidth requirements and latency in FL systems, particularly in edge computing scenarios. This approach enhances communication efficiency and opens new avenues for maintaining model quality under stringent bandwidth constraints. A significant advantage of using error-bounded lossy compression is that it can retain more information than the above methods. While compressors introduce noise, they do not erase information like sparsification or quantization at certain error boundaries. Developing error-bounded lossy compression strategies to target communication reduction for distributed learning systems is an open area of research. More work is needed to address this and to understand the nuances of controlled compression algorithms and how they affect model accuracy and performance.
@@ Acknowledgments
This research was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under contract DE-AC02-06CH11357, and supported by the National Science Foundation under Grant OAC-2003709, OAC-2104023, OAC-2311875, OAC-2330367, OAC-2311756, and OAC-2313122. The authors extend their appreciation to the Deanship of Scientific Research at University of Bisha, Saudi Arabia for funding this research work through the Promising Program under Grant Number (UB- Promising -40 - 1445).

### v3

@@ 6. Customized Compressors for Specific Applications or Use Cases
In this section, we comprehensively survey error-bounded lossy compressors specifically tailored to particular applications and use cases. We include these applications in our survey primarily because the associated compression methods explicitly support error bounding or preserve quantities of interest (QoIs), thus ensuring fidelity relevant to their respective domains.
Takeaway 4: Each use-case or application projects non-trivial requirements on posthoc analysis, compression ratio or performance, which often needs a codesign with corresponding domain experts.
@@ 6.1 Compression for Molecular Dynamics Simulations
Lossy compression has been widely considered a promising solution to reduce the data volume of MD simulations. There is ongoing research into lossy compression methods tailored for MD simulations. The HRTC method (Huwald et al., 2016) employs a strategy that represents trajectories as piecewise linear segments, coupled with quantization that is controlled for errors and a representation using variable-length integers. The PMC approach (Dvořák et al., 2020) leverages the information about atomic bonds within molecules to forecast the positions of atoms in each frame; however, this technique does not apply to simulations involving nonbonded interactions. Omeltchenko et al. (Omeltchenko et al., 2000) proposed a spatial compressor for MD datasets that includes three steps: (1) converting all floating-point values (both position and velocity) to integer numbers, (2) building a uniform oct-tree index according to the space-filling curve of the position fields, and (3) sorting the particles based on R-indices using a radix-similar sorting method in each block and encoding the difference in adjacent indices by variable-length encoding. Tao et al. (Tao et al., 2017b) improved Omeltchenko et al.’s method by sorting the particles based on a partial-radix sorting algorithm while preserving the same compression ratios by using SZ to compress the reordered coordinates instead of directly using the R-index. Essential dynamics (Meyer et al., 2006) is a powerful analysis tool for identifying the nature and relative importance of the essential deformation modes of a macromolecule from MD samplings. ED offers lossy compression by adopting principal component analysis (PCA) over the full trajectories of all particles. By comparison, Kumar et al. (Kumar et al., 2013) combined PCA and discrete cosine transform (DCT) to compress the full trajectories of the total MD dataset. Such full-trajectory-based compression methods, however, may be impractical for many of today’s large-scale MD simulations. Zhao et al. (Zhao et al., 2022) proposed an SZ2-based compressor, MDZ, which is equipped with spatial-clustering-based prediction and two-level temporal prediction. MDZ achieves high compression ratios particularly on MD simulations focusing on crystalline materials or featuring a continuous temporal domain.
@@ 6.3 Compression for Quantum Circuit Simulations
Wu et al. (Wu and et al., 2019) designed a Schrödinger algorithm-based full state vector simulation pipeline that integrates compression. MPI is used to parallelize the matrix multiplication required to apply a gate to a state vector. Each rank stores a set of compressed blocks that together compose a component of the overall state vector. Wu et al. explored multiple compressors to compress the state vector data, including SZ2.1 (A), SZ2.1 with complex type support (B), XOR leading-zero reduction coupled with bit-plane truncation and zstd (C), and reshuffling of real and imaginary parts together before performing C (D). The results indicate that solutions C and D achieve the highest compression ratios (30–90 for error bounds in the range [1E-5,1E-1]). When running a 61-qubit Grover’s search algorithm with this compressor, the memory requirement drops from 32 exabytes to 768 terabytes using 4,096 nodes. In all, their method can raise the number of qubits for a simulation by 2 to 16.
@@ 6.6 Compression for Topological Data Analysis
Topological data analysis is essential in abstracting, summarizing, and understanding scientific data in various applications, ranging from cosmology and combustion to Earth simulations and AI. Topological feature descriptors, or simply topological descriptors, provide robust capabilities for capturing, summarizing, and comparing features in scientific data. Most lossy compressors cannot preserve topological features, thus not guaranteeing topology preservation in decompressed data. Below is a review of customized lossy compressors in terms of topological data analysis from two aspects: scalar field topology and vector field topology.
@@ 6.7 Compression for Multi-resolution Data
Multi-resolution methods, such as the Adaptive Mesh Refinement (AMR) technique have been widely used in scientific simulations (Zhang et al., 2019a; Stone et al., 2020; Dubey and et al., 2022). AMR aims to reduce computational expenses while preserving the accuracy of simulation outcomes. Unlike traditional uniform mesh techniques that apply consistent resolution throughout the simulation space, AMR employs a dynamic approach. It selectively increases resolution in regions of interest, thereby optimizing computational resource usage and minimizing storage requirements. However, the space saved from using multi-resolution alone is often not enough. For instance, a multi-resolution dataset with $0.5\times 1024^{3}$ mesh points at the coarse level and $0.5\times 2048^{3}$ at the fine level could yield about 1 TB of data per snapshot. Consequently, conducting five simulations each with 200 snapshots require a total storage of 1 PB.
@@ 6.8 Compression for Seismic Imaging
Seismic imaging is a technique for determining the seismic properties of the Earth’s subsurface (Li and Qu, 2022). The technique is extensively utilized in earthquake imaging and resource exploration, including hydrocarbon and geothermal, by energy companies such as Saudi Aramco (Huang et al., 2023c). Reverse time migration (RTM) is a cutting-edge seismic imaging method since it can effectively analyze complex seismic structures (e.g., complex velocity focusing and steep (¿70) dips imaging), compared with traditional methods such as Kirchhoff and wave equation migration  (Farmer et al., 2009).
@@
Huang et al. (Huang et al., 2023c) proposed a hybrid lossy compression method called HyZ that combines blockwise regression (BR) and SZx (Yu et al., 2022) to improve the performance of RTM overall execution. Evaluation on 3,600 snapshots of the Overthrust model shows HyZ achieves a compression ratio of 12.31x and compression/decompression speeds of 10.69 GB/s and 12.45 GB/s, respectively. Integration of HyZ into an industrial parallel RTM code improves overall performance by 6.29–6.60x over the execution without compression techniques, outperforming second-tier compressors like SZ and ZFP by up to 2.23x. HyZ also demonstrates higher fidelity than BR in preserving the visualization quality of single snapshots and the final stacking image.
@@ 6.9 Compression for X-ray Light Source Data
For serial crystallography, two major approaches can be combined: non-hit rejection and ROIBIN-SZ (Underwood et al., 2023b). Like ptychography, the data present as 2D float-encoded integer data over a time dimension. Unlike ptychography, however, the data are substantially noisy, and there are features called Bragg spots or peaks that are key to the posthoc analysis. In order to achieve high compression ratios, a method called ROBIN-SZ was developed. ROIBIN-SZ uses the peak information used to perform non-hit rejection to losslessly preserve rectangular regions around the Bragg peaks, while using aggressive $2\times 2$ binning followed by SZ3 compression on the background. Preserving the background is critical because the peak-finding process is not infallible. There could be false-negative peaks to preserve in the dataset in order for the analysis process to complete.
@@ 6.10 Compression for Data Transfer over WAN
Recently, error-bounded lossy compression techniques have also been used to improve the data transfer performance over the wide area network (WAN). Ocelet (Liu et al., 2023a)—a lossy-compression-based data transfer accelerator developed for the Globus platform—is a typical example. The Ocelet framework is composed of 8 components/modules: user interface, FuncX service (fun, [n. d.]), Globus service (glo, [n. d.]), parallel executor, MPI call module, error-bounded lossy compression module, data loader/writer, and lossy compression quality estimation module. The parallel executor is used to launch the compression/decompression work in a parallel job. Globus manages the data transfer. FuncX service deals with remote orchestration. The user interface offers a graphical interface that helps users submit the tasks easily.
@@ 6.11 Compression for Boosting Communication in HPC Clusters
Researchers have been actively investigating the application of lossy compression to boost the performance of communication in high-performance clusters, focusing on two primary categories: point-to-point and collective communication.
@@ 6.12 Compression for Distributed Machine Learning & Federated Learning Systems
Recent years have witnessed the rapid evolution of deep learning models for getting high model accuracy, especially in the realm of large-scale models (Bommasani et al., 2022). Typical examples include large language foundation models (e.g., Palm and GPT-4 (Anil et al., 2023; OpenAI et al., 2023)), large-scale models in computer vision (e.g., VGGs, ResNets (Simonyan and Zisserman, 2015; He et al., 2016)), and life science (e.g., AlphaFolds (Jumper and et al., 2021)). Distributed systems, including the public/private cloud and HPC, provide strong support for training large-scale models based on large-volume training datasets to accelerate the training procedure  (Zhang et al., 2022; Zhang and Wang, 2021; Xu et al., 2024).
Communication between computing nodes during training has been recognized as the primary bottleneck in distributed training systems. During the training, gradients, model parameters, and activation data are transmitted across various nodes. As model sizes increase, the volume of this data escalates significantly, resulting in substantial communication overhead. To enhance training performance, it is essential to compress these data to reduce communication demands.
@@
Error-Bounded Lossy Compressors An emerging area of research is using error-bounded lossy compressors to reduce the size of model updates for distributed systems. Early approaches such as DeepSZ (Jin et al., 2019) focused on providing a framework for compression and storage of general DNN model architectures. This work showed that through compressing the weights of the model, high compression ratios could be achieved, leading to reported 50$\approx$ 50$\times$ 50 × compression on the weights. Other studies, such as the FedSZ framework (Wilkins et al., 2023), demonstrate how SZ-based compression can be integrated into federated learning (FL). By compressing local model updates before transmission, FedSZ effectively reduces the bandwidth requirements and latency in FL systems, particularly in edge computing scenarios. A significant advantage of using error-bounded lossy compression is that it can retain more information than the above methods. While compressors introduce noise, they do not erase information like sparsification or quantization at certain error boundaries. Developing error-bounded lossy compression strategies to target communication reduction for distributed learning systems is an open area of research.
6.13 Compression for Other Applications
In addition to the above-listed 12 representative scientific applications, there are more scientific applications that may generate masses of data to store and analyze potentially, including Plasma Physics Simulations (PPS) and Computational Fluid Dynamics (CFD). PPS, especially those employing Particle-in-Cell (PIC) methods commonly used in fusion energy research, can generate massive amounts of data (from tens to hundreds of terabytes per simulation). This is primarily due to the requirement to accurately model kinetic behaviors by tracking billions or even trillions of particle states across high-dimensional phase spaces (including positions, velocities, and electromagnetic fields). Recent research has shown that advanced compression approaches, including neural-network-based techniques (e.g., variational autoencoders) (Choi et al., 2021) and unsupervised machine-learning checkpoint strategies utilizing Gaussian Mixture Models (Chen et al., 2021a), hold great promise for alleviating storage and I/O challenges, facilitating efficient data handling, and enabling effective scientific analysis at extreme computing scales. As for CFD, it may also produce vast volumes of data to store and analyze, in that it typically requires very fine computational grids to accurately capture complex fluid phenomena (e.g., turbulence, shock waves, boundary layers) and involves quite a few physical variables (velocity components, pressure, density, temperature, turbulence quantities) at every grid point. Some compression methods have been developed for the compression of CFD data. BigWhoop (big, [n. d.]), for instance, is a compression library for numerical datasets. Lee et al. (Lee et al., 2024) also explored machine Learning techniques for data reduction of CFD applications.
@@ Acknowledgments
This research was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under contract DE-AC02-06CH11357 and DE-SC0024559, and supported by the National Science Foundation under Grant OAC-2104023, OAC-2311875, OAC-2311876, OAC-2311878, OAC-2330367, OAC-2311756, OAC-2311878, OAC-2313123, OAC-2344717, SHF-1943114, SHF-1910197, OAC-2313122 and NSF-2211538. The authors extend their appreciation to the Deanship of Graduate Studies and Scientific Research at University of Bisha for funding this research through the promising program under grant number (UB-Promising - 40 -1445).